Let F(x)=x^2-16 and G(x)= 1-x
Find (F-G)(-10)
(F-G)(-10)= simplify your answer

Answer by rothauserc(4718)

(F - G) (x) = x^2 - 16 - (1 - x)
:
(F - G) (x) = x^2 - 16 -1 + x
:
(F - G) (x) = x^2 +x - 17
:
(F - G) (-10) = (-10)^2 + (-10) + 17
:
(F - G) (-10) = 100 - 10 + 17
:
(F - G) (-10) = 107
: